### v1

@@ 4.1 Overall Workflow
Our voice-directed SAOP is a hierarchical multi-agent framework that automatically plans and executes the overall workflow (Fig. 1; Fig. 8). We divide the surgical video into 10-second clips because it is often longer than one hour, and each clip is designed to handle a single command. The workflow orchestrator agent first initializes the memory state, consisting of a local memory state for each clip and a global memory state to ensure consecutive information sharing across clips. The agent then leverages the LLM to automatically determine the next function to execute based on the available set of functions. The input prompt to the LLM includes instructions, function names and definitions, predifined workflow decision rules, the current status, and the output format requirements (Supplementary Note 4).
We categorize the functions into two types: workflow functions and task-specific agent functions. Workflow functions are for higher-level coordination that process and analyze voice commands through voice interaction, transcription, correction, validation, and reasoning for agent routing. Task-specific agent functions operate at a lower level to select the appropriate agent function and determine the state parameters. This hierarchical design improves modularity and maintainability, allowing the platform to flexibly incorporate new agents or functions without disrupting the overall workflow.
@@
In determining return points or termination, relying solely on function names and descriptions can lead to mispredictions. To mitigate this, workflow decision rules guide the overall process and define when to return and end. In addition, current status such as “No audio recorded”, “Last command invalid, need new input”, “IR Agent completed, workflow finished” provides contextual cues that help navigate branching points. The workflow decision rules and current status allow the orchestrator to make more accurate and consistent decisions across diverse scenarios.
@@
The LLM output should be in json format, containing each function name along with its associated probability for the next step. Incorporating probabilities allows the platform to provide a quantitative measure of confidence for decision-making. Since the LLM output may omit certain functions, the missing function completion rule includes all absent functions $M$ with their probabilities set to zero. Subsequently, the function selection rule chooses the next function $f_{next}$ with the highest probability over the complete function set. This hybrid approach combines the capabilities of LLMs with the probability-based decision rules, ensuring reliability and mathematical grounding.
4.2 Workflow Functions
4.2.1 Real-time Audio Function
@@
The real-time audio function manages voice interaction from capturing voice input from the local notebook to transmitting the audio file to a remote server as shown in Algorithm 1. As this function requires fast real-time processing on CPU, it integrates Silero-VAD (https://github.com/snakers4/silero-vad) for voice activity detection (VAD), Whisper-small model [44] via faster-whisper (https://github.com/SYSTRAN/faster-whisper/tree/master) for speech recognition and wake-word detection “davinci”, and Microsoft Edge TTS (https://github.com/rany2/edge-tts) for text-to-speech. To improve wake word detection and avoid false activation, we optimize Whisper-small decoding parameters by defining “davinci” as hotwords, setting the temperature to 0, beam size to $5$, no-speech threshold to $0.25$, compression ratio threshold to $1.2$, and log probability threshold to 1.0-1.0. The process starts with local notebook microphone streaming, where VAD detects speech activity, Whisper transcribes speech, identifies the wake word ”davinci”, records the surgeon’s command, and transmits to the remote server.
Algorithm 1 Real-time Audio Recording and Server Transmission Pipeline
Input: Audio stream from microphone, Silero-VAD model $M_{VAD}$, Whisper-small model $M_{W}$, Microsoft Edge TTS model $M_{TTS}$, keyword set $\mathcal{K}=\{\text{"davinci"},\text{"davin"},\text{"vinci"}\}$
Output: Recorded audio file transmitted to remote server
  1. 1.
    Initialization:
    • Load $M_{VAD}$ and $M_{W}$.
    • Play $M_{TTS}$(“Ready.”)
  2. 2.
    Keyword Detection (Wake-word Listening):
    • Open microphone stream $S$.
    • Continuously read chunks $x_{t}$ and compute timestamps $T=\text{VAD}(x_{t})$.
    • If speech detected by $M_{VAD}$, transcribe buffer via $M_{W}$.
    • If $\text{keyword}\in\mathcal{K}$ found, play $M_{TTS}$(“Please continue.”) and proceed.
  3. 3.
    Recording: Record speech until 4 seconds of silence.
  4. 4.
    Save and Transfer: Save as WAV, upload via SCP
4.2.2 Speech-to-Text (STT) Function
@@
The STT function employs the Whisper-medium model [44] integrated with faster-whisper for efficient inference. Comparing SpeechBrain model [45], the Whisper models, and Microsoft Phi-4-multimodal model [46], the Whisper-medium model offers a practical balance between speed and accuracy (Supplementary Note 5; Supplementary Fig. 4). The larger models generally achieve higher accuracy but incur slower inference. The Whisper-medium model is configured with a beam size of $8$, the language parameter set to “en”, a temperature of 0, a speech threshold of $0.3$, and a log prob threshold of 2.0-2.0 to enhance transcription accuracy for English speech while effectively filtering out non-speech segments and low-confidence outputs. Lastly, the transcribed command is updated in the memory state, and the overall procedure is shown in Algorithm 2.
Algorithm 2 Speech-to-Text (STT) Function
Input: Workflow state $\mathcal{S}$ containing audio path, STT model $M_{STT}$
Output: Transcription results (commands)
  1. 1.
    Initialization: Load audio from $\mathcal{S}[\text{audio\_path}]$.
  2. 2.
    Transcription:
    • Use $M_{STT}$ to transcribe the audio
    • Store each segment’s start time, end time, and text as stt_chunks.
  3. 3.
    Command Extraction:
    • Concatenate all transcribed texts to form commands.
    • If $\text{commands}=\emptyset$, set $\text{commands}=\text{“None”}$.
    • Append commands to $\mathcal{S}[\text{commands}]$.
  4. 4.
    Update workflow state: $\mathcal{S}\leftarrow\{\text{stt\_chunks},\ \text{commands}\}.$
4.2.3 Correct and Validate Function
@@
The correct and validate function employs Gemma3 [22] LLM to correct the transcribed command and determine its validity (Fig. 9a). The prompt contains the available task agents and their descriptions, instructions, command correction rules, validation rules, output format, memory state, and current transcribed command (Supplementary Note 6). The instruction first requests to translate the command into English and apply appropriate revisions. Since the STT model is unfamiliar with medical terminology such as axial, coronal, posterior, and CT, we provide text correction rules to handle such failure cases. For example, “Replace ‘city’ with ‘CT’” and “Replace COVID-related terms with ‘coronal’”. Ambiguity also arises with commands like zoom in/out, which may refer either to the IV agent or the AR agent. To resolve these issues, the prompt incorporates the memory state by providing the last three revised commands and associated agents. The LLM then outputs a json containing the corrected command and its validity, which is updated in the memory state. Valid commands are ready to proceed to the command reasoning function for further processing, whereas invalid ones should return to the real-time audio function for a new voice input. In addition, when the command is “None”, the platform automatically inherits the action from the previous video clip, and thus LLM revises as “Select {agent name}”.
Figure 9: The input and output of correct and validate function and command reasoning function. a The correct and validate predicts the corrected command and its validity. b The command reasoning prompt determines the most appropriate task-specific agent. Both functions leverage previous commands and agent history from global memory state, and step-by-step CoT reasoning.
4.2.4 Command Reasoning Function
@@
The command reasoning function takes the revised command as input and employs Gemma3 model to reason and select the most appropriate task agent (Fig. 9b). The prompt includes the available task agents with brief descriptions, instructions, memory state, and the current revised command (Supplementary Note 7). As in the correct and validate function, providing the last three revised commands and corresponding agents aids in resolving ambiguous commands. We apply Chain-of-Thought (CoT) prompting [23] for the reasoning process, which are helpful to analyze error cases and refine detailed agent descriptions according to them. The model finally returns a json object specifying the name of the selected agent, which is also updated in the memory state.
4.3 Task-specific Agent Functions
4.3.1 Information Retrieval (IR) Agent Functions
@@
The IR agent overlays the requested clinical data on the surgical video (Fig. 10; Supplementary Fig. 5). It comprises three functions: an action determiner (AD) function that interprets commands and decides the action and parameters, and two functions for showing or removing the overlay. The AD function leverages Gemma3 to infer the action uniquely mapped to an agent function and state parameters (Supplementary Note 8). We denote the command by $x\in\mathcal{X}$, the instruction by $i$, and a set of patient-data columns by $\mathcal{C}=\{c_{1},\dots,c_{m}\}$ with corresponding values $v=\{v_{1},\dots,v_{m}\}$. The LLM parameterized by $\theta$ produces an action distribution $\pi_{\theta}$ over $\{\text{SHOW},\text{HIDE}\}$ and a per-field inclusion probability $\phi_{\theta}$:
$\pi_{\theta}(a\mid x,i)=\Pr(A=a\mid x,i),$
@@
The agent selects the action $a^{\star}$ with the higher probability and predicts the indicator vector $y^{\star}\in\{0,1\}^{m}$ that specifies which fields to display, as formulated in Equation 6. To handle ambiguous expressions like physical information, the prompt guides their mapping to specific fields in $\mathcal{C}$, adjusting the $\phi_{\theta}$.
$a^{\star}=\arg\max_{a\in\{\text{SHOW},\text{HIDE}\}}\pi_{\theta}(a\mid x,i),$
@@
Figure 10: The overview of the IR agent. The LLM in the action determination function predicts the action and its state parameters, and the IR agent then executes either the show or remove function with inferred parameters.
@@
$\Psi=(y,s),\quad\begin{aligned} &y\in\{0,1\}^{m},\\ &s\in\text{final string}.\end{aligned}$ (7)
@@
For each selected column $c_{j}\in\mathcal{C}_{y^{\star}}$ with value $v_{j}$, a field-specific formatting function $r(c_{j},v_{j})$ transforms the raw data into a textual representation.
@@
This formatting function reflects physician feedback, such as combining pulmonary function test values in liters and percentages, merging relevant fields like sex and age, and omitting unnecessary units. The final patient information string is the result of concatenating all the formatted outputs with newline delimiters.
@@
4.3.2 Image Viewer (IV) Agent Functions
Figure 11: The overview of the IV agent. The action determination function predicts the action and state parameters, and the IV agent executes show, zoom in, zoom out, or remove function with inferred parameters.
The IV agent overlays CT DICOM images onto the surgical video, allowing surgeons to scroll through axial, coronal, and sagittal planes (Fig. 11; Supplementary Fig. 2). The AD function interprets commands to manipulate CT slices with four functions: displaying and moving multi-planar views, zooming into a large single view, zooming out to the small multi-planar views, and removing all CT views. We denote the command $x\in\mathcal{X}$, the instructions by $i$, and the state parameters by $\Psi$. The Gemma3 model parameterized by $\theta$ infers the action $A\in\mathcal{A},\;\mathcal{A}=${SHOW_MOVE, ZOOM_IN_MOVE, ZOOM_OUT, REMOVE} and updates the state parameters $\Psi$ (Supplementary Note 9). The decision process follows a joint policy that factorizes into an action selection policy and a state update policy:
$\pi_{\theta}(a,\Psi\mid x,i)$ $=\pi_{\theta}(a\mid x,i)\;\pi_{\theta}(\Psi\mid x,i,a),$
$a^{\star}$ $=\arg\max_{a\in\mathcal{A}}\ \pi_{\theta}(a\mid x,i),$
$\Psi^{\star}$ $=\arg\max_{\Psi}\ \pi_{\theta}(\Psi\mid x,i,a^{\star}),$ (12)
The policy first infers the action $a^{\star}$, and then predicts the updated state parameters $\Psi^{\star}$ conditioned on that action. The $\Psi$ denotes the current state and $\Psi^{\star}$ is the updated state after adjustments.
Specifically, the state parameters are defined as $\Psi=(p,\delta,v)$, where $p=(p^{\mathrm{axi}},p^{\mathrm{cor}},p^{\mathrm{sag}})$ denotes the slice positions, $\delta\in\{\text{none},\text{small\_views},\text{zoom\_view}\}$ is the display mode, and $v\in\{\text{axial},\text{coronal},\text{sagittal}\}$ is the main view plane. The slice positions are updated based on movement commands:
@@ 4.3.2 Image Viewer (IV) Agent Functions
where $\Delta p$ encodes positional adjustments such as “move right” or “sagittal plus 50” (increase sagittal index), “move down” or “axial minus 100” (decrease axial index), and “move forward” or “coronal plus 30” (increase coronal index). The display mode $\delta$ determines how CT slices appear on the surgical video. If $a^{\star}$ is SHOW_MOVE, $\delta$ is small_views, showing three CT views $I_{t}^{\mathrm{axi}},I_{t}^{\mathrm{cor}},I_{t}^{\mathrm{sag}}$ on the right side of the video, each with the updated positions $p^{\star}$. If $a^{\star}$ is ZOOM_IN_MOVE, $\delta$ is zoom_view, displaying a main CT view $I_{t}^{\mathrm{main}}$ in the center of the video. The main view $v$ is either explicitly specified in the command or implicitly inferred from the movement axis: $I_{t}^{\mathrm{main}}=\mathcal{V}(p^{\star},v)$. If $a^{\star}=\text{ZOOM\_OUT}$, $\delta$ is small_views, removing the main CT view and keeping the small views visible. If $a^{\star}=\text{REMOVE}$, $\delta$ is none, which removes all CT views.
@@ 4.3.3 Anatomy Rendering (AR) Agent Functions
The AR agent renders 3D anatomical models on the surgical video to enable interactive visualization (Fig. 12; Supplementary Fig. 3). The 3D models are reconstructed from CT images using TotalSegmentator (https://github.com/wasserth/TotalSegmentator) in the 3D Slicer (https://www.slicer.org/), software which generates segmented anatomical structures with corresponding labels. The AD function interprets commands and selects one of five actions: displaying static views, rotating the model, zooming in on specific structures, zooming out, and removing all models. Given a command $x\in\mathcal{X}$, instructions $i$, and the current state parameters $\Psi$, the Gemma3 LLM jointly infers the action $A\in\mathcal{A},\mathcal{A}=$ {STATIC, ROTATE, ZOOM_IN, ZOOM_OUT, REMOVE} and updates the state parameters $\Psi$ (Supplementary Note 10). The decision process follows the same joint policy formulation Equation  12 as in the IV agent. The key difference is that the actions, the state parameters, and the instructions to update them are all specific to the AR Agent. Specifically, the state parameters include active anatomical structures $\alpha$, viewpoint $v$, rotation mode $r$, target structure for zoom $\tau$, and zoom parameters $z$.
@@
The active anatomical structures $\alpha$ represent the list of visible structures. The physician selects the anatomical labels of interest, including left lower lobe, left upper lobe, right lower lobe, right middle lobe, right upper lobe, lung nodules, and the trachea/bronchia. The default setting activates only the lobe with nodules and all non-lobe structures. The LLM applies this default configuration when the command first invokes the AR agent, resets, or requests to show the surgical part. Subsequent commands that add or remove structures update $\alpha$, whereas commands that do not mention structure leave $\alpha$ unchanged.
@@
The viewpoint $v\in$ {anterior, posterior, left, right, superior, inferior, surgical} specifies the initial view of the 3D model. For example, commands such as “Show the anterior view” or “Show from the front” display the anterior side. In rotation commands, such as “Rotate the posterior view to the left”, the initial viewpoint $v$ is set to the posterior view. The default viewpoint is set to surgical view, which aligns to the perspective of a patient lying on the operating table, as illustrated in Figure 12.
@@
The rotation mode $r\in$ {static, left, right, up, down, horizontal, vertical} is the relative rotation direction from the initial viewpoint $v$. The static mode does not rotate the model, left/right/up/down modes rotate $30$ degrees toward the specified direction and returns to the original view, and the horizontal/vertical modes perform a continuous $360$ degrees rotation. The default setting is the static mode and clinicians can rotate the model in desired directions to inspect critical regions for resection.
@@
If $a^{\star}$ is ZOOM_IN or ZOOM_OUT, the LLM extracts the target structure $\tau$ from the command and updates the zoom parameters $z^{\star}=(z_{c}^{\star},z_{s}^{\star},z_{\ell}^{\star})$, where $z_{c}^{\star}$ is the center of $\tau$, $z_{s}^{\star}$ is the zoom scale, and $z_{\ell}^{\star}$ is the zoom depth level. The zoom in operation moves toward $z_{c}^{\star}$ and stores the zoom path in a history stack, which is later traversed in reverse order during the zoom out operation. The zoom scale $z_{s}$ has three values: start scale $z_{s}^{\text{start}}$, end scale $z_{s}^{\text{end}}$, and the actual current scale $z_{s}^{\text{current}}$. Specifically,
rotate/static: $\ z_{s}^{\text{start}}=z_{s}^{\text{end}}=z_{s}^{\text{current}},$
zoom in: $\ \begin{aligned} z_{s}^{\text{start}}&=z_{s}^{\text{current}},\quad z_{s}^{\text{end}}&=2\,z_{s}^{\text{current}}\end{aligned}$
zoom out: $\ \begin{aligned} z_{s}^{\text{start}}&=z_{s}^{\text{current}},\quad z_{s}^{\text{end}}&=\max\!\left(1.0,\ \tfrac{1}{2}z_{s}^{\text{current}}\right).\end{aligned}$ (15)
The scale starts at $z_{s}=1.0$, doubling with each zoom in and halving with each zoom out until the minimum scale of $1.0$. The zoom depth level $z_{\ell}$ increases for zoom in and decreases for zoom out.

### v3

@@ II-A Overall Workflow
Our VISA framework is a hierarchical MAS that automatically plans and executes the overall workflow as shown in Fig. 2. We divide the surgical video into 10-second clips, with each clip designed to handle a single command. The workflow orchestrator agent initializes the memory state for the target video clip and uses the LLM to select the next function to execute. The LLM prompt includes instructions, function names with definitions, predifined workflow decision rules, the current status, and the output format.
@@
In determining return points or termination, relying solely on function names and descriptions can lead to mispredictions. To mitigate this, workflow decision rules guide the overall process and define when to return and end. In addition, current status such as “No audio recorded”, “Last command invalid, need new input”, and “Agent completed, workflow finished” provide contextual cues that help navigate branching points. The workflow decision rules and current status allow the orchestrator to make more accurate and consistent decisions in various scenarios.
@@
The LLM output should be in json format, containing each function name along with its associated probability for the next step. Incorporating probabilities allows the system to provide a quantitative measure of confidence for decision-making. Since the LLM output may omit certain functions, the missing function completion rule includes all absent functions $M$ with their probabilities set to zero. Subsequently, the function selection rule chooses the next function $f_{next}$ with the highest probability over the complete function set. This hybrid approach combines the capabilities of LLMs with the probability-based decision rules, ensuring reliability and mathematical grounding.
II-B Workflow Functions
II-B1 Real-time Audio Function
@@
This function handles voice interaction by capturing audio on the edge device and transmitting it to a remote server. For fast real-time processing on CPU, it integrates Silero-VAD [^1] https://github.com/snakers4/silero-vad for voice activity detection (VAD), Whisper-small model [44] via faster-whisper [^2] https://github.com/SYSTRAN/faster-whisper for speech recognition and wake-word detection, and Microsoft Edge TTS [^3] https://github.com/rany2/edge-tts for text-to-speech (TTS). The process begins with microphone streaming, where VAD detects speech activity, Whisper transcribes the audio to identify the wake-word ”davinci”, records the command until four seconds of silence, and sends it to the remote server. Wake-word detection is further improved by optimizing Whisper-small decoding parameters, setting “davinci” as a hotword, the temperature to 0, beam size to $5$, no-speech threshold to $0.25$, compression ratio threshold to $1.2$, and log probability-threshold to 1.0-1.0.
II-B2 Speech-to-Text Function
@@
The STT function employs the Whisper-medium model [44] via faster-whisper for efficient inference. Among the SpeechBrain model [45], the Whisper models, and Microsoft Phi-4-multimodal model [46], the Whisper-medium model offers a practical balance between speed and accuracy. Testing on 35 randomly selected commands, larger models generally achieve higher accuracy but incur slower inference. The Whisper-medium model is configured with a beam size of $8$, the language set to “en”, a temperature of 0, a speech threshold of $0.3$, and a log prob threshold of 2.0-2.0 to enhance transcription accuracy for English speech while effectively filtering out non-speech segments and low-confidence outputs. Lastly, the transcribed command is updated in the memory state.
II-B3 Correct and Validate Function
@@
This function employs Gemma3 [22] LLM to correct the transcribed command and determine its validity. The prompt [^4] Prompts available at https://github.com/helena-lena/SAOP includes the available task agents with their descriptions, instructions, the current transcribed command, correction rules, validation rules, the output format, and the global memory state. Since the STT model is unfamiliar with medical terminology, we provide correction rules such as replacing “city” with “CT” or COVID-related terms with “coronal”. Ambiguity also arises with commands like “zoom in/out”, which may refer either to the IV agent or the AR agent. To resolve these problems, we provide the last three revised commands and associated agents from the global memory state. If the command is “None”, the LLM revises it to “Select {agent name}” by inheriting the agent from the previous memory state. The LLM outputs a json containing the corrected command and its validity, which is updated in the memory state. Valid commands proceed to the command reasoning function, while invalid ones return to the real-time audio function for a new voice input.
II-B4 Command Reasoning Function
@@
This function takes the revised command as input and employs Gemma3 model to reason and select the most appropriate task-specific agent. The prompt00footnotemark: 0 includes the available task agents with brief descriptions, instructions, the current revised command, and global memory state. Providing the last three revised commands and their corresponding agents from the global memory state helps the model resolve ambiguous commands. We also apply Chain-of-Thought (CoT) prompting [23] to the reasoning process to analyze the cause of failures and use these insights to refine agent descriptions. The model finally returns a json specifying the name of the selected agent, which is also updated in the memory state.
II-C Task-specific Agent Functions
II-C1 IR Agent Functions
@@
The IR agent overlays the requested clinical data on the surgical video and consists of three functions: an action determiner (AD) function that interprets commands and decides the action and parameters, and two functions for showing or removing the overlay as shown in Fig. 3. The AD function leverages Gemma3 to infer the action uniquely mapped to an agent function and state parameters00footnotemark: 0. We denote the command by $x\in\mathcal{X}$, the instruction by $i$, and a set of patient data columns by $\mathcal{C}=\{c_{1},\dots,c_{m}\}$ with corresponding values $v=\{v_{1},\dots,v_{m}\}$. The LLM parameterized by $\theta$ produces an action distribution $\pi_{\theta}$ over $\mathcal{A}=\{\text{SHOW},\text{HIDE}\}$ as in (1) and a per-field inclusion probability $\phi_{\theta}$ as in (2):
$\pi_{\theta}(a\mid x,i)=\Pr(\mathcal{A}=a\mid x,i),$ (1)
@@
The agent selects the action $a^{\star}$ with the higher probability and predicts the indicator vector $y^{\star}\in\{0,1\}^{m}$ that specifies which fields to display, as in (4). To handle ambiguous expressions like physical information, the prompt guides their mapping to specific fields in $\mathcal{C}$, adjusting the $\phi_{\theta}$.
$a^{\star}=\arg\max_{a\in\{\text{SHOW},\text{HIDE}\}}\pi_{\theta}(a\mid x,i),$ (3)
@@
$\Psi=(y,s),\quad\begin{aligned} &y\in\{0,1\}^{m},&s\in\text{final string}.\end{aligned}$ (5)
@@
For each selected column $c_{j}\in\mathcal{C}_{y^{\star}}$ with value $v_{j}$, a field-specific formatting function $r(c_{j},v_{j})$ transforms the raw data into a textual representation as in (7).
@@
This formatting function reflects physician feedback, such as combining pulmonary function test values in liters and percentages, merging relevant fields like sex and age, and omitting unnecessary units. The final patient information string is the result of concatenating all the formatted outputs with newline delimiters as in (8).
@@
Figure 3: The overview of the IR agent, which executes either the show or remove function with inferred state parameters.
Figure 4: The overview of the IV agent, which executes show, zoom in, zoom out, or remove function with inferred state parameters.
II-C2 IV Agent Functions
The IV agent overlays CT DICOM images onto the surgical video, allowing surgeons to scroll through axial, coronal, and sagittal planes as shown in Fig. 4. The AD function interprets commands and selects one of four actions: displaying and moving multi-planar views, zooming into a large single view, zooming out to the small multi-planar views, and removing all CT views00footnotemark: 0. Given the command $x\in\mathcal{X}$ and the instructions $i$, the Gemma3 model parameterized by $\theta$ infers the action $a\in\mathcal{A}$ and updates the state parameters $\Psi$ following the joint policy:
$\pi_{\theta}(a,\Psi\mid x,i)$ $=\pi_{\theta}(a\mid x,i)\;\pi_{\theta}(\Psi\mid x,i,a),$ (10)
$a^{\star}$ $=\arg\max_{a\in\mathcal{A}}\ \pi_{\theta}(a\mid x,i),$ (11)
$\Psi^{\star}$ $=\arg\max_{\Psi}\ \pi_{\theta}(\Psi\mid x,i,a^{\star}).$ (12)
The policy first infers the action $a^{\star}$, and then predicts the updated state parameters $\Psi^{\star}$ conditioned on that action. The $\Psi$ denotes the current state and $\Psi^{\star}$ is the updated state.
Specifically, the state parameters are defined as $\Psi=(p,\delta,v)$, where $p=(p^{\mathrm{axi}},p^{\mathrm{cor}},p^{\mathrm{sag}})$ is the slice positions, $\delta\in\{\text{none},\text{small\_views},\text{zoom\_view}\}$ is the display mode, and $v\in\{\text{axial},\text{coronal},\text{sagittal}\}$ is the main view plane. The slice positions are updated based on movement commands:
@@ II-C2 IV Agent Functions
where $\Delta p$ encodes positional adjustments such as “move right” or “sagittal plus 50” (increase sagittal), “move down” or “axial minus 100” (decrease axial), and “move forward” or “coronal plus 30” (increase coronal). The display mode $\delta$ determines how CT slices appear on the surgical video. If $a^{\star}$ is SHOW_MOVE, $\delta$ is small_views, showing three CT views $I_{t}^{\mathrm{axi}},I_{t}^{\mathrm{cor}},I_{t}^{\mathrm{sag}}$ on the right side of the video, each with the updated positions $p^{\star}$. If $a^{\star}$ is ZOOM_IN_MOVE, $\delta$ is zoom_view, displaying a main CT view $I_{t}^{\mathrm{main}}$ in the center of the video. The main view $v$ is either explicitly specified in the command or implicitly inferred from the movement axis: $I_{t}^{\mathrm{main}}=\mathcal{V}(p^{\star},v)$. If $a^{\star}=\text{ZOOM\_OUT}$, $\delta$ is small_views, removing the large main CT view. If $a^{\star}=\text{REMOVE}$, $\delta$ is none, which removes all CT views.
@@ II-C3 AR Agent Functions
The AR agent renders 3D anatomical models onto the surgical video for interactive visualization as shown in Fig. 5. The 3D models are reconstructed from CT images using TotalSegmentator [^5] https://github.com/wasserth/TotalSegmentator in the 3D Slicer [^6] https://www.slicer.org/, which generates anatomical structures with corresponding labels. The AD function interprets commands and selects one of five actions: displaying static views, rotating the model, zooming in on specific structures, zooming out, and removing all models00footnotemark: 0. The Gemma3 LLM jointly infers the action $a$ and updates the state parameters $\Psi$ following the same joint policy formulation as (12) in the IV agent. The difference is that $a$, $\Psi$, and $i$ are all specific to the AR Agent. The state parameters $\Psi$ include active anatomical structures $\alpha$, viewpoint $v$, rotation mode $r$, target structure for zoom $\tau$, and zoom parameters $z$.
@@
The active structures $\alpha$ represent the list of visible anatomical labels. The physician selects the labels of interest, including left lower lobe, left upper lobe, right lower lobe, right middle lobe, right upper lobe, lung nodules, and the trachea/bronchia. The default setting activates the lobe containing the nodules together with all non-lobe structures, and subsequent commands add or remove structures accordingly.
@@
The viewpoint $v\in$ {anterior, posterior, left, right, superior, inferior, surgical} defines the initial view of the 3D model. The default viewpoint is set to surgical view, which aligns to the perspective of a patient lying on the operating table as shown in Figure 5. For example, commands like “Show the anterior view” or “Show from the front” display the anterior side, while rotation commands such as “Rotate the posterior view to the left” set the $v$ to the posterior view.
@@
The rotation mode $r\in$ {static, left, right, up, down, horizontal, vertical} is the relative rotation direction from the initial viewpoint $v$, with the static mode used by default. The static mode leaves the model fixed, the left/right/up/down modes rotate by $30$ degrees toward the specified direction and return, and the horizontal/vertical modes perform a continuous $360$ degrees rotation.
@@
If $a^{\star}$ is zoom in/out, the LLM extracts the target structure $\tau$ from the command and updates the zoom parameters $z^{\star}=(z_{c}^{\star},z_{s}^{\star},z_{\ell}^{\star})$, where $z_{c}^{\star}$ is the center of $\tau$, $z_{s}^{\star}$ is the zoom scale, and $z_{\ell}^{\star}$ is the zoom depth level. The zoom in moves toward $z_{c}^{\star}$ and stores the zoom path in a history stack, which is later traversed in reverse order during the zoom out. The zoom scale starts at $z_{s}=1.0$, doubles when zooming in, and halves when zooming out until reaching the minimum scale of $1.0$, while the zoom depth level increases or decreases accordingly.